OpenStudy (anonymous):

Bobby is bringing his collection of dimes and quarters to the bank to deposit into his savings account. If d represents the number of dimes in Bobby’s collection, q represents the number of quarters, and the collection is worth $63.75, which of the following equations models the value of Bobby’s collection?

OpenStudy (anonymous):

??

OpenStudy (anonymous):

A. 63.75 = d + q B. 63.75 = 25d + 10q C. 63.75 = 10d + 25q D. 63.75 = 0.10d + 0.25q

OpenStudy (anonymous):

since dime is 10 cents and quarter is 25 cents. So if you pug in the values, what would satisfy?

OpenStudy (anonymous):

A?

OpenStudy (anonymous):

if it is only d+q, that will be 10+25, so 35 cents. so nope! give it another tryyyy :)

OpenStudy (anonymous):

is it D because those are in cents

OpenStudy (anonymous):

hmm if you make $63.75 to cents, what would it be?

OpenStudy (anonymous):

6375

OpenStudy (anonymous):

YASSSS! great job! now which equation would fit to 6375?

OpenStudy (anonymous):

hey are you sure that you typed them right???

OpenStudy (anonymous):

yah I'm pretty sure but none of the equations seem to match

OpenStudy (anonymous):

huh... thats weired.. hold up

OpenStudy (anonymous):

what is this from??

OpenStudy (anonymous):

the online practice course i am taking.

OpenStudy (anonymous):

ohh do you think thee could have been a mistake when someone was making this?

OpenStudy (anonymous):

there could be let me email the instructor, thanks for trying to help though!!

OpenStudy (anonymous):

welcome!!!